# Rate-induced biosphere collapse in the Daisyworld model

### v1

@@ II The Daisyworld model
The Daisyworld model is specified by the two coupled ordinary differential equations:
@@
Here $\alpha_{w},\alpha_{w}$ are the fraction of the planet’s surface covered by white and black daisies, respectively. $\alpha_{g}=1-\alpha_{b}-\alpha_{w}$ is the fraction of uncovered ground, while $\beta(T)$ is the growth rate and $\gamma$ the death rate of the daisies. It is assumed that there are no nutrient limitations, and that the entire surface of the planet is in principle suitable for hosting daisies. The growth rates of both daisy populations depend quadratically on temperature, with some optimum temperature $T_{\rm opt}$:
@@
The effective emission temperature of the planet, $T_{e}$, is given by the radiative balance:
$\sigma T_{e}^{4}=SL(1-A),$ (4)
where $S$ is a radiative flux in units of W/m2, and $L$ is a dimensionless parameter. For simplicity the planet is assumed to have no greenhouse gases in its atmosphere, and the factor of $1/4$ that arises in energy balance equations for rapidly rotating spherical planets is neglected. $A$ is the albedo (reflectivity) of the planet: it is given by
@@
where $A_{w}$ and $A_{b}$ are the albedos of the white and black daisies and $A_{g}$ is the albedo of uncovered ground. Finally, it is assumed that, due to the different albedos, each of the three kinds of surface (white daisy, black daisy, uncovered) has its own local temperature:
@@
with $i$ replaced by $b,w$, or $g$ as appropriate. $q$ is a positive constant quantifying the horizontal heat transfer across regions with different albedos. The outgoing radiative fluxes can thus be summed to recover the total outgoing flux:
@@
To facilitate comparisons with previous studies, the parameter values we use are almost all taken directly from Watson and Lovelock (1983). We use $\gamma=0.3$, $k=0.003265$, $T_{\rm opt}=295.5$ K, $S=917$ W/m2, $A_{w}=0.75$, $A_{g}=0.5$, and $A_{b}=0.25$. It is worth noting that the death rate $\gamma$ defines a natural biotic turnover timescale $\tau_{b}=1/\gamma\approx 3.3$. The heat transfer coefficient $q$ cannot be larger than $SL/\sigma$, as this implies a transfer of heat against the temperature gradient. Watson and Lovelock (1983) used fixed values of $q$ less than $0.2SL/\sigma$. In this study we vary $L$ rapidly within the range of about $0.6-1.6$; for simplicity, we set $q=0.1S/\sigma$.
@@
Figure 1: (a) Equilibrium states in the Daisyworld model: $\alpha_{b}$ is the fraction of the planet covered by black daisies, $\alpha_{w}$ is the fraction of the planet covered by white daisies, and $L$ is the dimensionless luminosity (incoming sunlight). There are six physically relevant equilibrium states, which we label as $e_{0}$ through to $e_{5}$. $e_{0}$ is a dead planet, $e_{1}$ and $e_{2}$ are states in which there are only white daisies, $e_{3}$ and $e_{4}$ are states in which there are only black daisies, and $e_{5}$ is a "coexistence" state, in which there are both black and white daisies. Solid lines indicate stable equilibria, and dashed lines unstable equilibria. (b) Phase portraits for different values of the luminosity $L$. Filled circles are stable equilibria, and unfilled circles are unstable equilibria. Of particular interests are the stable manifolds of certain saddle points: these are highlighted in blue for easy future reference.
The phase portraits in Figure 1 further show some example phase space trajectories. Of particular interest are the stable manifolds of certain saddle points ($e_{4}$ for smaller values of $L$, and $e_{1}$ for larger values of $L$): these are highlighted in blue for easy reference later on.
@@
Figure 2: Effective emission temperature $T_{e}$ as a function of varying luminosity $L$ for the different equilibrium states of the model. Notably, when daisies are present, $T_{e}$ tends to be maintained at values that are much more conducive to the survival of daisies. For example, for larger values of $L$ (ca. 1.2-1.5) white daisies increasingly dominate, increasing the albedo of the planet and reducing overall temperature relative to a dead planet. Conversely, for smaller values of $L$ black daisies increasingly dominate, decreasing albedo and increasing overall temperature relative to a dead planet. This is the widely-discussed "self-regulation" of Daisyworld’s biosphereWatson and Lovelock (1983); Wood et al. (2008).
@@ III Bifurcation-induced collapse
in which $\Lambda(rt)$ is an input parameter that varies on a much slower time scale $r\,t$. Here, $r$ is a small dimensionless parameter which we refer to as the rate parameter.
@@
exhibits a "dangerous" bifurcationThompson, Stewart, and Ueda (1994) at some $\lambda=\lambda^{*}$, then the nonautonomous system in Eq. 8 may tip when $\Lambda(rt)$ crosses $\lambda^{*}$. Specifically, this will occur if the system state is initially in a stable equilibrium which loses stability as a result of the bifurcation, and $\Lambda(rt)$ remains beyond $\lambda^{*}$ for a sufficiently long time interval. As discussed in the Introduction, bifurcation-induced tipping is responsible for fixed (i.e. time-invariant) thresholds ($\lambda^{*}$in this case).
@@
In the Daisyworld model, increasing $L$ to a large enough value, or decreasing it to a small enough value, will lead to bifurcation-induced tipping towards extinction. This is already apparent from the left-hand panel in Figure 1, and we demonstrate this in more detail in Figure 3, where we consider two scenarios.
@@
Figure 3: Bifurcation-induced tipping towards extinction in the Daisyworld model. In both (a) and (b) we initialise the system at a stable "living" equilibrium (blue dot), and then gradually change the luminosity $L$. In panel (a), there is initially a large population of white daisies; however, eventually the equilibrium undergoes a saddle-node bifurcation (red dot), the only remaining stable equilibrium is that of a dead planet, and the biosphere collapses. In panel (b) we demonstrate the same phenomenon starting from the stable state in which there are only black daisies, and decreasing $L$. In each case, dashed lines denote unstable equilibrium states.
@@
In the first scenario, Figure 3 (a), we initialise the system at fairly high luminosity $(L=1.4)$, from the stable equilibrium in which only white daisies exist. Then, as luminosity is increased, eventually the stable equilibrium undergoes a saddle-node bifurcation, and only the "dead planet" state is now stable. As a result, the system inevitably evolves towards the dead planet state, and the daisies go extinct. Physically, the planet’s surface temperature was initially maintained by the white daisies far below where it would otherwise have been. However, the effect of increased temperature on daisy growth rates (Eq. 3) eventually overwhelms the stabilising effect of the daisy population on temperature, and the daisy population abruptly collapses.
@@
In the second scenario, Figure 3 (b), we initialise the system at fairly low luminosity $(L=0.68)$, from a state in which only black daisies exist, We now decrease the luminosity until the equilibrium disappears through a saddle node bifurcation, with similar consequences as in the first case. Initially the planet’s temperature was maintained by the black daisies far above where it would otherwise have been. But again, the effect of temperature on daisy growth rates eventually overwhelms the stabilising effect of the daisy population on temperature, and the daisy population collapses.
IV Rate-induced collapse via basin instability
@@
Next, and more notably, we show that Daisyworld’s biosphere can undergo rate-induced tipping towards extinction. This is demonstrated in Figure 4. We initialise the system at the coexistence equilibrium with some smaller $L$ value we call $L_{\rm min}$, and consider the time-dependent forcings $L(rt)$ which ultimately increase $L$ to some larger value $L_{\rm max}$. The coexistence equilibrium exists and is stable in the entire range $L_{\rm min}\leq L\leq L_{\rm max}$. The forcings take the specific form
$L(rt)=L_{\min}+\frac{\Delta L}{2}(\tanh(rt)+1)$ (10)
with the rate parameter $r$, which is not necessary small here, controlling how quickly the transition between $L_{\rm min}$ and $L_{\rm max}$ occurs.
In Figure 4 we show the result of such a perturbation for two different values of $r$: one fast (red), and one slow (green). In both cases, the change is fast enough that the system state fails to track the moving equilibrium ($e_{5}(L)$). In the slow case, the system is able to eventually reach this equilibrium ($e_{5}(L_{\rm max})$): the biosphere has survived the change in $L$. However, when $L$ changes more quickly (red), the biosphere collapses, leaving a dead planet. The collapse occurs even though a stable biosphere could in principle exist across the entire range $L_{\rm min}\leq L\leq L_{\rm max}$. This is rate-induced tipping.
@@
Figure 4: Rate-induced tipping towards extinction in the Daisyworld model. We initialise the system at the coexistence equilibrium $(e_{5})$ with a value of $L=L_{\rm min}$. Then, we ramp $L$ towards some new value $L_{\rm max}$, at a speed determined by a dimensionless rate parameter $r$. Panel (b) shows time series of $L(rt)$, while panels (c) and (d) show time series of $\alpha_{w}$ and $\alpha_{b}$. When the change is slow enough (green, $r=0.5$), the biosphere survives; when it is too fast (red, $r=1$), the biosphere collapses. This occurs because the system can enter the basin of attraction of the "dead planet" state, $e_{0}$, if the change in $L$ is fast enough. The threshold for entering this basin of attraction is the stable manifold of the saddle equilibrium $e_{1}$ (see also Figure 5 and panel f of Figure 1). In the space of $\alpha_{w}$,$\alpha_{b}$, and $L$, the threshold becomes a two dimensional surface $W^{s}(e_{1}(L))$: trajectories which cross it are precisely those which tip. This is shown in panel (a).
To understand why this occurs, we can consider Figure 5, which shows certain equilibrium states and basins of attraction for different values of $L$. At $L=L_{\max}$, the phase space is partitioned into two separate basins of attraction by the stable manifold of the saddle equilibrium $e_{1}$ (blue curve): the basin of the stable coexistence equilibrium $e_{5}(L_{\max})$ (light blue), and the basin of the dead planet equilibrium $e_{0}(L_{\max})$ (white). For fixed luminosity $L=L_{\max}$, trajectories starting in the former basin converge to the stable coexistence equilibrium $e_{5}$, while trajectories starting in the latter basin converge to the "dead planet" equilibrium $e_{0}$. Note that the stable coexistence equilibrium at $L_{\min}$, $e_{5}(L_{\min})$ (grey point) lies in the basin of attraction of $e_{0}(L_{\max})$. As such, the equilibrium $e_{5}$ is called basin unstableO’Keeffe and Wieczorek (2020). Figure 5 (b) shows the two basins at $L=L_{BI}$, which is the value of the parameter where the equilibrium $e_{5}(L_{\min})$ intersects the boundary of the basins. We refer to $L_{BI}$ as the boundary of the basin instability region of equilibrium $e_{5}(L_{\min})$.
@@
To explain rate-induced tipping, when $L(rt)$ changes from $L_{\min}$ to $L_{\max}$, we consider two extreme cases. First, when $r\to 0$, $L(rt)$ changes extremely slowly. Thus, the system can keep track of the moving equilibrium $e_{5}(L(rt))$. Second, when $r\to\infty$, $L(rt)$ changes instantaneously from $L_{\min}$ to $L_{\max}$. Hence, the state of the system, although it starts at $e_{5}(L_{\min})$, ends up in the basin of $e_{0}(L_{\max})$. Therefore, the system tips. By continuity, there is a critical value of the rate parameter $r_{c}$ such that the system exhibits tipping for any $r>r_{c}$ and tracks for $r<r_{c}$. If $r=r_{c}$ the system converges to the saddle point $e_{1}(L_{\max}))$Wieczorek, Xie, and Ashwin (2023). (We note that transition behaviour in dynamical systems due to crossing stable manifolds of saddle points has long been considered, see e.g. FitzHugh (1955))
@@
Figure 5: Rate-induced tipping in the Daisyworld model (e.g. as seen in Figure 4) is a consequence of basin instability. Panel (a) shows that, at $L=L_{\max}$ (the upper end of the perturbation in Figure 4) the phase space is partitioned into two basins of attraction, separated by the stable manifold of the saddle point $e_{1}(L_{\max})$. The stable coexistence equilibrium $e_{5}(L_{\min})$ lies within the basin of attraction of $e_{0}(L_{\max})$. Thus, if the system is initialised at $e_{5}(L_{\min})$ with $L=L_{\min}$, and then $L$ is instantaneously changed to $L_{\max}$, the system tips towards $e_{0}$ (the "dead planet" state). The equilibrium $e_{5}$ is thus basin unstableO’Keeffe and Wieczorek (2020). Panel (b) shows the two basins at $L=L_{BI}$, which is where $e_{5}(L_{\min})$ intersects the boundary of the basins. We refer to this as the boundary of the basin instability region of equilibrium $e_{5}(L_{\min})$.
@@
We can further characterise the kinds of perturbations that trigger rate-induced tipping. We continue to initialise the system at the stable coexistence equilibrium $e_{5}(L_{\min})$, and perturb it using the forcing function in Eq. 10, varying both the rate parameter $r$ and the total change in $L$, $\Delta L=L_{\rm max}-L_{\rm min}$. Our results are shown in Figure 6, with the two examples from Figure 4 again highlighted. For fast enough rates $r$, the $\Delta L$ required to tip the system converges to a constant value, which represents the $\Delta L$ value required to instantaneously move the system state across the basin instability boundary. For slower perturbations, the system is able to remain in the original basin of attraction, and is thus always able to recover to the original (but now shifted) coexistence state.
@@
Figure 6: Tipping diagram for the Daisyworld model, starting from the same coexistence equilibrium $e_{5}$ with luminosity $L=L_{\rm min}$. Panels (b), (c), and (d) again show time series of $\alpha_{b}$,$\alpha_{w}$, and $L(rt)$, for the same two examples as in Figure 4. Panel (a) shows, for different combinations of rate $r$ and total luminosity change $\Delta L=L_{\rm max}-L_{\rm min}$, which perturbations result in tipping, and which ones do not. For large rates $r$, the critical $\Delta L$ converges to a fixed value, which represents the instantaneous $\Delta L$ required to move the system across the basin threshold.
@@
One further conclusion from this diagram is that, on Daisyworld, for a general perturbation in $L$, basin instability (i.e. rate-induced tipping) is often "easier" to initiate than bifurcation-induced tipping. To see this, we begin by looking at Figure 4, and noting that we start from a value of $L_{\rm min}=0.8$. For an increase in $L$ which is monotonic but whose time dependence is otherwise arbitrary, the furthest that $L$ could possibly increase while Daisyworld’s biosphere survives is until the saddle-node bifurcation at $L\simeq 1.55$: see Fig. 1 panel a, and Fig 3, panel a. (We note that, by this point, the stable living equilibrium is no longer the coexistence state $e_{5}$, but instead the "white daisies only" state $e_{2}$.) Thus, the point at which $\Delta L$ must necessarily trigger extinction is at $\Delta L\simeq 1.55-0.8=0.75$. But, as Figure 4 shows, perturbations of sufficient speed can trigger extinction for much smaller values of $\Delta L$. On Daisyworld, rates of change can play as important a role in determining catastrophic outcomes as the magnitude of the change.

### v2

@@ II The Daisyworld model
The Daisyworld model is specified by the two coupled ordinary differential equations
@@
Here $\alpha_{w},\alpha_{w}$ are the fraction of the planet’s surface covered by white and black daisies, respectively. $\alpha_{g}=1-\alpha_{b}-\alpha_{w}$ is the fraction of uncovered ground, while $\beta(T)$ is the growth rate and $\gamma$ the death rate of the daisies. It is assumed that there are no nutrient limitations, and that the entire surface of the planet is in principle suitable for hosting daisies. The growth rates of both daisy populations depend quadratically on temperature, with some optimum temperature $T_{\rm opt}$.
@@
Here, $k$ is a parameter which determines how sensitive the growth rate is to temperature.
The effective emission temperature of the planet, $T_{e}$, is given by the radiative balance
$\sigma T_{e}^{4}=SL(1-A).$ (4)
$\sigma T_{e}^{4}$ is the radiation emitted by the planet, and $SL(1-A)$ is the radiation received from Daisyworld’s host star. $S$ is a radiative flux in units of W m-2. We follow Watson and Lovelock (1983) in referring to $L$ as a dimensionless luminosity: scaling the star’s luminosity by some factor is equivalent to scaling insolation (the actual radiation arriving at the planet’s surface) by the same factor.
For simplicity the planet is assumed to have no greenhouse gases in its atmosphere, and is assumed to be flat, such that the factor of $1/4$ that arises in energy balance equations for rapidly rotating spherical planets is neglected. $A$ is the albedo (reflectivity) of the planet: it is given by
@@
where $A_{w}$ and $A_{b}$ are the albedos of the white and black daisies and $A_{g}$ is the albedo of uncovered ground. Finally, it is assumed that, due to the different albedos, each of the three kinds of surface (white daisy, black daisy, uncovered) has its own local temperature
@@
with $i$ replaced by $b,w$, or $g$ as appropriate. $q$ is a positive constant quantifying the horizontal heat transfer across regions with different albedos. The outgoing radiative fluxes can thus be summed to recover the total outgoing flux.
@@
To facilitate comparisons with previous studies, our parameter values are all the same as those of Watson and Lovelock (1983), except for the heat transfer coefficient $q$ (see below). We use $\gamma=0.3$ (unit time)-1, $k=0.003265$ (unit time)-1 K-2, $T_{\rm opt}=295.5$ K, $S=917$ W m-2, $A_{w}=0.75$, $A_{g}=0.5$, and $A_{b}=0.25$. $\sigma$ is the Stefan-Boltzmann constant, which we take as 5.67 $\times$ 10-8 W m-2 K-4. The heat transfer coefficient $q$ cannot be larger than $SL/\sigma$, as this implies a transfer of heat against the temperature gradient. Watson and Lovelock (1983) used fixed values of $q$ less than $0.2SL/\sigma$. In this study we vary $L$ rapidly within the range of about $0.6-1.6$; for simplicity, we set $q=0.1S/\sigma$.
As in the original modelWatson and Lovelock (1983) the unit of time is left unspecified. (This causes no inconsistencies with the usage of W m-2 for fluxes due to how the equations are framed.) It is worth noting that the death rate $\gamma$ defines a natural biotic turnover timescale (i.e. timescale of biosphere renewal) $\tau_{b}=1/\gamma\approx 3.3$.
@@
Figure 1: (a) Equilibrium states in the Daisyworld model: $\alpha_{b}$ is the fraction of the planet covered by black daisies, $\alpha_{w}$ is the fraction of the planet covered by white daisies, and $L$ is the dimensionless luminosity (incoming sunlight). There are six physically relevant equilibrium states, which we label as $e_{0}$ through to $e_{5}$. $e_{0}$ is a dead planet, $e_{1}$ and $e_{2}$ are states in which there are only white daisies, $e_{3}$ and $e_{4}$ are states in which there are only black daisies, and $e_{5}$ is a "coexistence" state, in which there are both black and white daisies. Solid lines indicate stable equilibria, and dashed lines unstable equilibria. (b-g) Phase portraits for different values of the luminosity $L$. Filled circles are stable equilibria, and unfilled circles are unstable equilibria. Of particular interests are the stable manifolds of certain saddle points: these are highlighted in orange for easy future reference.
@@
The phase portraits in Figure 1 further show some example phase space trajectories. Of particular interest are the stable manifolds of certain saddle points ($e_{4}$ for smaller values of $L$, and $e_{1}$ for larger values of $L$): these are highlighted in orange for easy reference later on.
@@
Figure 2: Effective emission temperature $T_{e}$ as a function of varying luminosity $L$ for the different equilibrium states of the model. Solid lines are stable equilibria, dashed lines are unstable equilibria, and filled circles are bifurcations. Notably, when daisies are present, $T_{e}$ tends to be maintained at values that are much more conducive to the survival of daisies. For example, for larger values of $L$ (ca. 1.2-1.5) white daisies increasingly dominate, increasing the albedo of the planet and reducing overall temperature relative to a dead planet. Conversely, for smaller values of $L$ black daisies increasingly dominate, decreasing albedo and increasing overall temperature relative to a dead planet. This is the widely-discussed "self-regulation" of Daisyworld’s biosphereWatson and Lovelock (1983); Wood et al. (2008).
@@ III Bifurcation-induced collapse
in which $\Lambda(rt)$ is an input parameter that varies on a much slower time scale $r\,t$. Here, $r$ is a small parameter whose unit is 1/(unit time). In this setting, the parameter $r$ quantifies the rate of change of the input parameter $\Lambda$, and is often referred to as the rate parameter Ashwin, Perryman, and Wieczorek (2017); Ritchie et al. (2023).
@@
exhibits a "dangerous" bifurcationThompson, Stewart, and Ueda (1994) at some $\lambda=\lambda^{*}$, then the nonautonomous system in Eq. 8 may tip when $\Lambda(rt)$ crosses $\lambda^{*}$. Specifically, this will occur if the system state is initially in a stable equilibrium which loses stability as a result of the bifurcation, and $\Lambda(rt)$ remains beyond $\lambda^{*}$ for a sufficiently long time interval. As discussed in the Introduction, bifurcation-induced tipping is responsible for fixed (i.e. time-independent) thresholds ($\lambda^{*}$in this case). In this case, as long as $\Lambda(rt)$ crosses $\lambda^{*}$, the precise value of $r$ does not affect whether or not tipping occurs.
@@
In the Daisyworld model, increasing $L$ to a large enough value, or decreasing it to a small enough value, will lead to bifurcation-induced tipping towards extinction. This is already apparent from the left-hand panel in Figure 1, and we demonstrate this in more detail in Figure 3. Here and throughout the rest of the paper, we implement time-varying forcings in luminosity $L$ according to
$L(rt)=L_{\rm start}+\frac{\Delta L}{2}(\tanh(rt)+1),$ (10)
where $\Delta L=L_{\rm end}-L_{\rm start}$, and the rate parameter $r$ controls how quickly the transition between $L_{\rm end}$ and $L_{\rm start}$ occurs. Unless specified otherwise, we will use values of $L_{\rm end}>L_{\rm start}$.
@@
Figure 3: Bifurcation-induced tipping towards extinction in the Daisyworld model. In both (a) and (b) we initialise the system at a stable "living" equilibrium (black dot), and then gradually change the luminosity $L(rt)$. In panel (a), there is initially a large population of white daisies; however, eventually the equilibrium undergoes a saddle-node bifurcation (magenta dot), the only remaining stable equilibrium is that of a dead planet, and the biosphere collapses. In panel (b) we demonstrate the same phenomenon starting from the stable state in which there are only black daisies, and decreasing $L$. In each case, dashed lines denote unstable equilibrium states. Panels (c), (d) show time series of the luminosity $L(rt)$ with $r=0.01$, corresponding to (a) and (b), respectively. Panels (e) and (f) show time series of $\alpha_{w}$ and $\alpha_{b}$ corresponding to (a) and (b), respectively.
@@
In Figure 3 we consider two scenarios. In the first scenario, shown in Figure 3 (a), (c), and (e), we initialise the system at fairly high luminosity $(L=1.4)$, from the stable equilibrium in which only white daisies exist. Then, as luminosity is increased, eventually the stable equilibrium undergoes a saddle-node bifurcation, and only the "dead planet" state is now stable. As a result, the system inevitably evolves towards the dead planet state, and the daisies go extinct. Physically, the planet’s surface temperature was initially maintained by the white daisies far below where it would otherwise have been. However, the effect of increased temperature on daisy growth rates (Eq. 3) eventually overwhelms the stabilising effect of the daisy population on temperature, and the daisy population abruptly collapses.
@@
In the second scenario, shown in Figure 3 (b), (d), and (f), we initialise the system at fairly low luminosity $(L=0.68)$, from a state in which only black daisies exist, We now decrease the luminosity (in this particular case $L_{\rm end}<L_{\rm start}$) until the equilibrium disappears through a saddle node bifurcation, with similar consequences as in the first case. Initially the planet’s temperature was maintained by the black daisies far above where it would otherwise have been. But again, the effect of temperature on daisy growth rates eventually overwhelms the stabilising effect of the daisy population on temperature, and the daisy population collapses.
There is no bifurcation-induced tipping starting directly from the coexistence equilibrium state, because the only bifurcations this can undergo are transcritical bifurcations to the single-species equilibria (see Figure 1). However, we note that if the system is initialised at the coexistence equilibrium state for any given value of $L$, and $L$ is slowly increased or decreased, for $|\Delta L|$ large enough the system will eventually undergo bifurcation-induced tipping: it is just that the system will pass through a single-species equilibrium first.
IV Rate-induced collapse and basin instability
@@
Next, and more notably, we show that Daisyworld’s biosphere can undergo rate-induced tipping towards extinction. This is demonstrated in Figure 4. We initialise the system at the coexistence equilibrium with some smaller $L$ value $L_{\rm start}$, and consider the time-dependent forcings $L(rt)$ (Eq. 10) which ultimately increase $L$ to some larger value $L_{\rm end}$. The coexistence equilibrium exists and is stable in the entire range $L_{\rm start}\leq L\leq L_{\rm end}$.
@@
Figure 4: Rate-induced tipping towards extinction in the Daisyworld model. We initialise the system at the coexistence equilibrium $(e_{5})$ with a value of $L=L_{\rm start}$. Then, we ramp $L$ towards some new value $L_{\rm end}$, at a speed determined by a dimensionless rate parameter $r$. Panel (b) shows time series of $L(rt)$, while panels (c) and (d) show time series of $\alpha_{w}$ and $\alpha_{b}$. When the change is slow enough (blue, $r=0.5$), the biosphere survives; when it is too fast (red, $r=1$), the biosphere collapses. This occurs because the system can enter the basin of attraction of the "dead planet" state, $e_{0}$, if the change in $L$ is fast enough. The threshold for entering this basin of attraction is the stable manifold of the saddle equilibrium $e_{1}$ (see also Figure 5 and panel f of Figure 1). In the space of $\alpha_{w}$,$\alpha_{b}$, and $L$, the threshold becomes a two dimensional surface $W^{s}(e_{1}(L))$: trajectories which cross it are precisely those which tip. This is shown in panel (a), in which the red curve crosses the surface and the blue curve does not.
In Figure 4 we show the result of such a perturbation for two different values of $r$: one fast (red), and one slow (blue). In both cases, the change is fast enough that the system state fails to track the moving equilibrium ($e_{5}(L)$). In the slow case, the system is able to eventually reach this equilibrium ($e_{5}(L_{\rm end})$): the biosphere has survived the change in $L$. However, when $L$ changes more quickly (red), the biosphere collapses, leaving a dead planet. The collapse occurs even though a stable biosphere could in principle exist across the entire range $L_{\rm start}\leq L\leq L_{\rm end}$. (i.e. the coexistence equilibrium is stable for all $L_{\rm start}\leq L\leq L_{\rm end}$). This is rate-induced tipping.
To understand rate-induced tipping, we consider Figure 5, which shows the coexistence stable equilibrium $e_{5}(L)$, the dead planet equilibrium $e_{0}(L)$, and their basins of attraction at different values of the parameter $L$, for the autonomous Daisyworld model.
As mentioned before, the coexistence equilibrium $e_{5}(L)$ is stable for luminosity $L$ in $[L_{\rm start},L_{\rm end}]$, and so is the dead planet equilibrium $e_{0}(L)$. This means that for all $L$ in $[L_{\rm start},L_{\rm end}]$ the phase space is partitioned into two basins of attraction. These are the basins of attraction of $e_{5}(L)$ and of $e_{0}(L)$. To simplify the discussion, we write
$B(e_{5},L),$
to denote the basin of attraction of $e_{5}$ at the parameter value $L$, and
$B(e_{0},L),$
to denote the basin of attraction of $e_{0}$ at the parameter value $L$. The boundary between these two basins is the stable manifold of the saddle equilibrium $e_{1}(L)$ (orange curve).
In Figure 5 we plot $B(e_{5},L)$ (light orange) and $B(e_{0},L)$ (white) for two different luminosity values $L=L_{\rm end}$ and $L=L_{BI}$, Additionally, we show the stable coexistence equilibrium $e_{5}(L_{\rm start})$ as a grey point, which represents the starting point for the R-tipping trajectories in Figure 4.
In panel (a) for $L=L_{\rm end}$, the equilibrium $e_{5}(L_{\rm start})$ lies outside the basin of attraction of $e_{5}(L_{\rm end})$, i.e.
$e_{5}(L_{\rm start})\not\in B(e_{5},L_{\rm end}).$
In such a case, the equilibrium $e_{5}(L_{\rm start})$ is said to be basin unstable on the parameter path $[L_{\rm start},L_{\rm end}]$ O’Keeffe and Wieczorek (2020).
Figure 5 (b) shows the two basins at $L=L_{BI}$, which is the value of the parameter where the equilibrium $e_{5}(L_{\rm start})$ intersects the boundary of the basins. We refer to $L_{BI}$ as the boundary of the basin instability region of equilibrium $e_{5}(L_{\rm start})$.
@@
To explain rate-induced tipping, when $L(rt)$ changes from $L_{\rm start}$ to $L_{\rm end}$, we consider two extreme cases. First, when $r\to 0$, $L(rt)$ changes extremely slowly. Thus, the system can keep track of the moving equilibrium $e_{5}(L(rt))$. Second, when $r\to\infty$, $L(rt)$ changes instantaneously from $L_{\rm start}$ to $L_{\rm end}$. Hence, the state of the system, although it starts at $e_{5}(L_{\rm start})$, ends up in the basin of $e_{0}(L_{\rm end})$. Therefore, the system tips. By continuity, there is a critical value of the rate parameter $r_{c}$ such that the system exhibits tipping for any $r>r_{c}$ and tracks for $r<r_{c}$. If $r=r_{c}$ the system converges to the saddle point $e_{1}(L_{\rm end})$Wieczorek, Xie, and Ashwin (2023). (We note that transition behaviour in dynamical systems due to crossing stable manifolds of saddle points has long been considered, see e.g. FitzHugh (1955)).
One can consider an augmented (compactified) system where $L(rt)$ can be incorporated as an additional differential equation. Such an augmented system will exhibit a codimension-one heteroclinic bifurcation for $r=r_{c}$ Wieczorek, Xie, and Ashwin (2023). For sufficiently simple models the heteroclinic orbit, and hence $r_{c}$, can be computed in closed form Ritchie and Sieber (2016) or approximated using perturbation methods. To highlight rate-induced biosphere collapse in the Daisyworld model, we find it sufficient to provide a numerical approximation of the critical rate parameter $r_{c}$.
@@
Figure 5: Rate-induced tipping in the Daisyworld model (e.g. as seen in Figure 4) is a consequence of basin instability. Panel (a) shows that, at $L=L_{\rm end}$ (the upper end of the perturbation in Figure 4) the phase space is partitioned into two basins of attraction, separated by the stable manifold of the saddle point $e_{1}(L_{\rm end})$. The stable coexistence equilibrium $e_{5}(L_{\rm start})$ lies within the basin of attraction of $e_{0}(L_{\rm end})$. Thus, if the system is initialised at $e_{5}(L_{\rm start})$ with $L=L_{\rm start}$, and then $L$ is instantaneously changed to $L_{\rm end}$, the system tips towards $e_{0}$ (the "dead planet" state). The equilibrium $e_{5}$ is thus basin unstableO’Keeffe and Wieczorek (2020). Panel (b) shows the two basins at $L=L_{BI}$, which is where $e_{5}(L_{\rm start})$ intersects the boundary of the basins. We refer to this as the boundary of the basin instability region of equilibrium $e_{5}(L_{\rm start})$.
@@
We can further characterise the kinds of perturbations that trigger rate-induced tipping. We continue to initialise the system at the stable coexistence equilibrium $e_{5}(L_{\rm start})$, and perturb it using the forcing function in Eq. 10, varying both the rate parameter $r$ and the total change in $L$, $\Delta L=L_{\rm end}-L_{\rm start}$. Our results are shown in Figure 6, with the two examples from Figure 4 again highlighted. For fast enough rates $r$, the $\Delta L$ required to tip the system converges to a constant value, which represents the $\Delta L$ value required to instantaneously move the system state across the basin instability boundary. For slower perturbations, the system is able to remain in the original basin of attraction, and is thus always able to recover to the original (but now shifted) coexistence state.
@@
Figure 6: Tipping diagram for the Daisyworld model, starting from the same coexistence equilibrium $e_{5}$ with luminosity $L=L_{\rm start}$. Panels (b), (c), and (d) again show time series of $\alpha_{b}$,$\alpha_{w}$, and $L(rt)$, for the same two examples as in Figure 4. Panel (a) shows, for different combinations of rate $r$ and total luminosity change $\Delta L=L_{\rm end}-L_{\rm start}$, which perturbations result in tipping, and which ones do not. For large rates $r$, the critical $\Delta L$ converges to a fixed value, which represents the instantaneous $\Delta L$ required to move the system across the basin threshold.
@@
One further conclusion from this diagram is that, on Daisyworld, for a general perturbation in $L$, basin instability (i.e. rate-induced tipping) is often "easier" to initiate than bifurcation-induced tipping. To see this, we begin by looking at Figure 4, and noting that we start from a value of $L_{\rm start}=0.8$. For an increase in $L$ which is monotonic but whose time dependence is otherwise arbitrary, the furthest that $L$ could possibly increase while Daisyworld’s biosphere survives is until the saddle-node bifurcation at $L\simeq 1.55$: see Fig. 1 panel a, and Fig 3, panel a. (We note that, by this point, the stable living equilibrium is no longer the coexistence state $e_{5}$, but instead the "white daisies only" state $e_{2}$.) Thus, the point at which $\Delta L$ must necessarily trigger extinction is at $\Delta L\simeq 1.55-0.8=0.75$. But, as Figure 4 shows, perturbations of sufficient speed can trigger extinction for much smaller values of $\Delta L$. On Daisyworld, rates of change can play as important a role in determining catastrophic outcomes as the magnitude of the change.